You have some chocolate bars which are 55% cocoa, and some other chocolate bars that are 25% cocoa. If you melt them down and mix them together, and end up with chocolate that is 49% cocoa, how many if each bar was used?
Answer by stanbon(75887)

You have some chocolate bars which are 55% cocoa, and some other chocolate bars that are 25% cocoa. If you melt them down and mix them together, and end up with chocolate that is 49% cocoa, how many if each bar was used?
--------
Equation:
cocoa + cocoa = cocoa
0.55A + 0.25B = 0.48(A+B)
------------------------------
55A + 25B = 48A + 48B
7A = 23B
-----
One possible answer:
23 of the 55% bars
7 of the 25% bars
=======================
Cheers,
Stan H.